# HetSwarm: Cooperative Navigation of Heterogeneous Swarm in Dynamic and Dense Environments through Impedance-based Guidance

### v1

@@ I Introduction
While heterogeneous may improve both the precision and time of logistics, their navigation suffers from the additional density of the swarm morphology. This paper introduces a HetSwarm setup that combines a drone and a ground robot in a leader-follower configuration, with the drone acting as the leader while the ground robot follows them through impedance link [7] connectivity.
@@
Furthermore, the heterogeneous setup significantly enhances efficiency for logistics tasks. A drone has the ability to carry only a limited payload due to its size and battery constraints whereas the ground robot can carry heavier payloads, thereby improving overall performance. Additionally, if the drone’s battery is depleted, it can land on the ground robot for recharging, ensuring uninterrupted operation.
@@ II Related Works
Batinovic et al. [8] utilized the APF method to address the path planning challenge for aerial robots operating in an unknown environment, focusing on ensuring safe trajectory execution and avoiding intricate obstacles using a LiDAR sensor. In a similar vein, Yu et al. [9] proposed an innovative distributed control algorithm that integrates the APF technique within a virtual leader formation scheme, coupled with a switching communication network. Malopolski et al. [10] presented an autonomous mobile robot for transport tasks in warehouses; a drive mechanism was proposed for surface and rail navigation and an elevator for vertical movement. However, the lack of aerial capabilities limits its ability to access hard-to-reach areas and navigate dense environments, reducing adaptability in multi-level spaces. Zhura et al. [11] investigated the impact of UAV in heterogeneous mapping and the navigation of a quadrupled robot. Sales et al. [12] developed a highly-scalable and low-cost multirobot system for inventory management composed of pairs with a micro-UAV and a ground mobile robot. Castro et al. [13] proposed the strategy to assist the cooperation of a heterogeneous robot team that involves two UGVs and one UAV. The robots operate in a partially known dynamic environment where they exchange information among themselves and perform their task of aerial and ground inspections.
@@
Khan et al. [14] proposed a leader-follower approach using an APF path planner and impedance controller for a multi-drone homogeneous system, allowing agents to plan the path and navigate to the target in an unknown yet static environment. However, this approach could not work in dynamic environment scenarios. Additionally, this research lacks the ability to operate over extended periods due to the limited battery life and flight time of the drones. To overcome the individual limitations of aerial and ground robots, heterogeneous swarm systems have been used. Chen et al. [15] target search and navigation and design a heterogeneous robot system consisting of a UAV and a UGV for search and rescue missions in unknown environments.
@@
The HetSwarm system, inspired by the previous work of SwarmPath [14] and SwarmGear [16], introduces a new agile and safe path for heterogeneous systems in dynamic and cluttered environments.
@@ III-A System Overview
Figure 2: Figure shows the overall system architecture and the pipeline of HetSwarm.
@@ III-B Artificial Potential Fields for Global Path Generation
In order for the leader drone to navigate efficiently around the obstacles while setting the path toward the goal, we applied the APF planning algorithm [17].The algorithm generates a virtual force that attracts the drone towards the target and a repulsive force that pushes it away from obstacles. An optimized trajectory is generated using the combination of these two forces. The equations for APF planner are as follows [14]:
$F_{\text{total}}=F_{\text{attraction}}+F_{\text{repulsion}}$ (1)
@@
$F_{\text{repulsion}}(d_{\text{o}})$ =$=\begin{cases}0&\text{if }d_{\text{o}}>d_{\text{safe}}\\ k_{\text{rep}}\cdot\left(\frac{1}{d_{\text{o}}}-\frac{1}{d_{\text{safe}}}\right)&\text{if }d_{\text{o}}\leq d_{\text{safe}}\end{cases}$
where $d_{\text{g}}$ and $d_{\text{o}}$ are the distances from drone to goal and to obstacle, respectively, $k_{\text{att}}$ and $k_{\text{rep}}$ are the attraction and repulsion coefficients, respectively.
@@ III-C1 Connectivity of Mobile Robot with Drone
In this configuration, the follower ground robot position is coupled with the leader-drone APF trajectory, which serves as the leading trajectory, through a mass-spring-damper system. Thereby, creating virtual impedance links among the drone and the ground robot. These impedance links provide a smooth connection among the agents. The links are established using a second-order differential equation of mass-spring-damper, which is given as:
$m\Delta\ddot{x}+d\Delta\dot{x}+k\Delta x=F_{\text{ext}}(t)$ (2)
where $\Delta x$ is the difference between the current and desired mobile robot position and $F_{\text{ext}}(t)$ is the virtual external force applied as an input from the leader drone. $m$ is the virtual mass of a link, $d$ is the damping coefficient of the virtual damper, and $k$ is the virtual spring constant.
@@ III-C2 Connectivity of Mobile Robot with Ground Obstacles
The APF trajectory generated by the leader drone does not account for smaller obstacles. To avoid collisions between the mobile robot and obstacles, we enabled the ground robot to establish additional links with the obstacles while simultaneously disconnecting from the drone, ensuring a collision-free path:
$\Delta x_{robot,n}=k_{impF}\cdot r_{imp}$ (3)
where $r_{imp}$ is the radius of the local deflection region around the obstacle, $k_{impF}$ is the force coefficient adjusted to the ground robot’s average velocity, and $n$ is the number of mobile robots.
@@ IV-A Experimental Setup in GymPybullet Environment
A custom simulation environment has been developed in the Gym PyBullet environment 3, incorporating the dynamics of both a drone and a mobile robot. The environment allows the visualization of multiple obstacles, and the obstacles can be moved dynamically. A custom PID controller has been implemented for the mobile robot that helps it to follow the path by adjusting its velocity and steering angles, while the drone’s PID controller ensures stable and controlled drone flight. This setup creates realistic interactions among the drone and mobile robot in dynamic environments, making it suitable for testing heterogeneous robots in real-world-like conditions. The drone utilizes the APF planner to generate and adjust its path in real-time; simultaneously, the mobile robot maintains the connectivity with the drone and ground obstacles using impedance links. To evaluate the adaptability and robustness of the system, experiments were performed on multiple cases. In a dynamic environment, three cases were performed from sparse to dense scenarios; cases were run thirty times to check the algorithm’s robustness and repeatability of performance. Similarly, in static environments, experiments were run under varying conditions to do a detailed analysis.
Figure 3: Experimental setup in Gym Pybullet environment showing the trajectory of drone and mobile robot under dense environment.
@@ IV-B1 Results in static environment
Furthermore, the mobile robot avoids obstacles encountered along its path, leading to impedance deflections that are also illustrated in the figure. The green circles represent ground-level obstacles, and the graph demonstrates that the drone’s trajectory remains straight, passing over these obstacles while the mobile robot effectively navigates around them. Fig. 4a, Fig. 4b, and Fig. 4c represent three cases with the same starting and goal positions of the agent, demonstrating how the algorithm effectively handles both sparse and dense scenarios. Furthermore, by changing the initial and target positions, different path planning scenarios can be tested, as shown in Fig. 4d.
@@ IV-B2 Results in dynamic environment
Real-time simulation was also performed on dynamic environments involving the continuous movement of one ground obstacle and one obstacle that was deflected by both the drone and mobile robot. Fig. 5 shows the tests that were performed in two main cases with sparse (Fig. 5a) and dense (Fig. 5b) environments. The change of path can be observed in the region where the movement of the obstacle takes place. The same behavior is observed in a cluttered environment, shown in Fig.5c and Fig.5d. The results revealed the deficient adaptable behavior of heterogeneous swarm under multiple environmental conditions.
Figure 5: Simulation examples of dynamic obstacle avoidance in sparse (a,b) and dense (c,d) environments. Dashed purple lines are the obstacle trajectories. Black dots are the obstacles for both robots, green dots are the ground obstacles deflected only by the mobile robot. Gray and light green-colored circles show the safe deflection regions for both agents.
@@ IV-B5 Time efficiency in task completion for multiple agents
To check the robustness and system adaptability under varying conditions, we conducted experiments on two different environments, shown in Fig. 4c, and Fig. 4d, with varying starting and goal positions, both in a static and dynamic environment. Table II shows how the trajectory lengths and mission time change under varying environmental conditions.
@@ IV-B6 Success Rate
Success Rate=(Successful ExperimentsTotal $\text{Success Rate}=\left(\frac{\text{Successful Experiments}}{\text{Total Experiments}}\right)\times 100$ (4)
@@ IV-C Comparison with Conflict-Based Search (CBS) Algorithm
1 0.73 0.25 0.75 0.38
2 0.72 0.25 0.75 0.37
With similar environmental configurations, the CBS trajectory in Case 1, where there are fewer obstacles, shows that agents tend to move around obstacles, resulting in a longer path. In contrast, APF agents navigate within obstacles, leading to more efficient paths. However, in Case 2, where the environment is more densely packed with obstacles, CBS agents ignore the inner regions and move around obstacles to complete the trajectory. On the other hand, APF agents move between obstacles, encountering multiple deflections, which increase their path length. Despite this, APF excels in terms of minimum obstacle distance; agents maintain a very close distance to obstacles while still avoiding collisions. In comparison, CBS agents maintain a minimum distance of 72 cm, reflecting a more cautious but less flexible path.
@@ IV-D Conclusion and Future Work
The research presents a heterogeneous swarm system comprising of a drone and the mobile robot, designed to collaborate in dynamic environments for logistics and warehouse management tasks. This research works on a leader-follower approach where a drone, being a leader, uses an APF path planner to navigate to goal positions while avoiding obstacles. Additionally, the mobile robot follows the drone using impedance links that connect the drone with the mobile robot. Moreover, the mobile robot develops additional impedance links with the ground obstacles to ensure smooth navigation toward the goal.
The system also provides efficient navigation of the swarm in dynamic and cluttered environments, thereby enhancing the adaptability and performance of the system in real-time scenarios. By combining the two agents, the system proves to have the complementary strengths of both a drone and a mobile robot. Multiple experiments were performed on this system, achieving a success rate of 90%. Additionally, a small deviation of about 45 cm was observed in the mobile robot’s path, particularly near the ground obstacles region, indicating the robot is effectively avoiding them.
The current approach was later compared with the Conflict-Based Search (CBS) algorithm, which reveals that our approach enables agents to safely yet optimally navigate within 25 cm of obstacles, whereas CBS maintains a minimum distance of 73 cm.

### v2

@@ I Introduction
While heterogeneous systems may improve both the precision and time of logistics, their navigation suffers from the additional density of the swarm morphology. This paper introduces a HetSwarm setup that combines a drone and a ground robot in a leader-follower configuration, with the drone acting as the leader while the ground robot follows them through impedance link [7] connectivity.
@@
Furthermore, the heterogeneous setup significantly enhances efficiency for logistics tasks. A drone has the ability to carry only a limited payload due to its size and battery constraints, whereas the ground robot can carry heavier payloads, thereby improving overall performance. Additionally, if the drone’s battery is depleted, it can land on the ground robot for recharging, ensuring uninterrupted operation.
@@ II Related Works
Controlling a large swarm of UAVs imposes a complex task. Several recent works propose to introduce humans in the loop for complex decision-making, e.g., Abdi et al. [8] with EMG-based gesture control and Khen et al. [9] proposing gesture recognition and machine learning. However, most of the systems rely on autonomous drones, with paths of vehicles being planned through various routing approaches. Batinovic et al. [10] utilized the APF method to address the path planning challenge for aerial robots operating in an unknown environment, focusing on ensuring safe trajectory execution and avoiding intricate obstacles using a LiDAR sensor. In a similar vein, Yu et al. [11] proposed an innovative distributed control algorithm that integrates the APF technique within a virtual leader formation scheme, coupled with a switching communication network. Malopolski et al. [12] presented an autonomous mobile robot for transport tasks in warehouses; a drive mechanism was proposed for surface and rail navigation and an elevator for vertical movement. However, the lack of aerial capabilities limits its ability to access hard-to-reach areas and navigate dense environments, reducing adaptability in multi-level spaces. Cooperative UAV and AGV teams previously were extensively explored for rescue missions; for example, Salas et al. [13] proposed a strategy based on the exploration of an unknown environment through the collaboration of the UAV equipped with a monocular camera and the differential-drive AGV. Zhura et al. [14] investigated the impact of UAVs in heterogeneous mapping and the navigation of a quadruped robot. Sales et al. [15] developed a highly scalable and low-cost multirobot system for inventory management composed of pairs with a micro-UAV and a ground mobile robot. Hajkarim et al. [16] developed a new mission planning optimization method based on reinforcement learning and multi-agent rollout policy optimization for coverage missions involving uncrewed aerial systems and ground vehicles to minimize the mission planning time and heterogeneous route length. Castro et al. [17] proposed the strategy to assist the cooperation of a heterogeneous robot team that involves two AGVs and one UAV. The robots operate in a partially known dynamic environment where they exchange information among themselves and perform their task of aerial and ground inspections. The decentralized control barrier function was proposed by Bhatia et al. [18] as a solution for distributed global connectivity maintenance for a multi-agent system.
@@
Tsykunov et al. [19] first suggested the concept of SwarmTouch that utilized virtual impedance links introduced by Hogan [7] for control of the homogeneous swarm of drones. Fedoseev et al. [20] further explored this concept by analyzing different impedance link topologies in the swarm formation. Khan et al. [21] proposed a leader-follower approach using an APF path planner and impedance controller for a multi-drone homogeneous system, allowing agents to plan the path and navigate to the target in an unknown yet static environment. However, the above-mentioned approaches could not work in a dynamic environment. Additionally, the ability to operate over extended periods due to limited battery life was not previously addressed. To overcome the individual limitations of aerial and ground robots, heterogeneous swarm systems have been used. Darush et al. [22] implemented links in a heterogeneous swarm as a formation control method between a leader octocopter and a follower swarm of micro-drones, potentially able to dock to the leader drone [23]. Chen et al. [24] developed a heterogeneous robot system for target search and navigation, consisting of a UAV and an AGV for rescue missions in unknown environments.
@@
The HetSwarm system, inspired by the previous work of SwarmPath and SwarmGear, introduces a new agile and safe path planner for heterogeneous systems in dynamic and cluttered environments, relying on guiding UAVs and heterogeneous impedance links to preserve formation control.
@@ III-A System Overview
Figure 2: System architecture and the pipeline of HetSwarm for global and local collision avoidance.
@@ III-B Artificial Potential Fields for Global Path Generation
In order for the leader drone to navigate efficiently around the obstacles while setting the path toward the goal, we applied the APF planning algorithm [25]. The algorithm generates a virtual force that attracts the drone towards the target and a repulsive force that pushes it away from obstacles. An optimized trajectory is generated using the combination of these two forces. The equations for the APF planner are as follows [21]:
$F_{\text{total}}=F_{\text{attraction}}+F_{\text{repulsion}},$ (1)
@@
$F_{\text{repulsion}}(d_{\text{o}})$ $=\begin{cases}0&\text{if }d_{\text{o}}>d_{\text{safe}}\\ k_{\text{rep}}\cdot\left(\frac{1}{d_{\text{o}}}-\frac{1}{d_{\text{safe}}}\right)&\text{if }d_{\text{o}}\leq d_{\text{safe}},\end{cases}$
where $d_{\text{g}}$ and $d_{\text{o}}$ are the distances from the drone to the goal and to the obstacle, respectively, $k_{\text{att}}$ and $k_{\text{rep}}$ are the attraction and repulsion coefficients, respectively.
@@ III-C1 Connectivity of Mobile Robot with Drone
In this configuration, the follower ground robot position is coupled with the leader drone APF trajectory, which serves as the guiding trajectory through a mass-spring-damper system and creates a virtual impedance link between the drone and the ground robot. These impedance links provide a smooth connection among the agents. The links are established using a second-order differential equation of mass-spring-damper, which is given as in [22]:
$m\Delta\ddot{x}+d\Delta\dot{x}+k\Delta x=F_{\text{ext}}(t),$ (2)
where $\Delta x$ is the difference between the current and desired mobile robot position and $F_{\text{ext}}(t)$ is the virtual external force applied as an input from the leader drone, $m$ is the virtual mass of a link, $d$ is the damping coefficient of the virtual damper, and $k$ is the virtual spring constant.
@@ III-C2 Connectivity of Mobile Robot with Ground Obstacles
The APF trajectory generated by the leader drone does not account for smaller obstacles. To avoid collisions between the mobile robot and obstacles, we enabled the ground robot to establish additional links with the low-height obstacles while simultaneously disconnecting from the drone, ensuring a collision-free path:
$\Delta x_{robot,n}=k_{impF}\cdot r_{imp},$ (3)
where $r_{imp}$ is the radius of the local deflection region around the obstacle, $k_{impF}$ is the force coefficient adjusted to the ground robot’s average velocity, and $n$ is the number of mobile robots. This equation ensures a collision-free trajectory by applying a repulsive displacement proportional to the obstacle’s influence radius $r_{imp}$, redirecting the robot away from potential collisions. The coefficient $k_{\text{impF}}$ adjusts the strength of this deflection based on the robot’s velocity, ensuring timely and stable avoidance.
@@ IV-A Experimental Setup in Gym Pybullet Environment
A custom simulation environment has been developed in the Gym PyBullet environment 3, incorporating the dynamics of both a drone and a mobile robot. The environment allows the visualization of multiple obstacles, and the obstacles can be moved dynamically. A custom PID controller has been implemented for the mobile robot that helps it to follow the path by adjusting its velocity and steering angles, while the drone’s PID controller ensures stable and controlled drone flight. This setup creates realistic interactions between the drone and mobile robot in dynamic environments, making it suitable for testing heterogeneous robots in real-world-like conditions. The drone utilizes the APF planner to generate and adjust its path in real-time; simultaneously, the mobile robot maintains the connectivity with the drone and ground obstacles using impedance links. To evaluate the adaptability and robustness of the system, experiments were performed on multiple cases. In a dynamic environment, three cases were performed from sparse to dense scenarios; cases were run thirty times to check the algorithm’s robustness and repeatability of performance. Similarly, in static environments, experiments were run under varying conditions to do a detailed analysis.
Figure 3: Experimental setup in the Gym Pybullet environment showing the trajectory of the drone and mobile robot under a dense environment.
@@ IV-B1 Results in static environment
Furthermore, the mobile robot avoids obstacles encountered along its path, leading to impedance deflections that are also illustrated in the figure. The green circles represent ground-level obstacles, and the graph demonstrates that the drone’s trajectory remains straight, passing over these obstacles while the mobile robot effectively navigates around them. Fig.4a-c represent three cases with the same starting and goal positions of the agent, demonstrating how the algorithm effectively handles both sparse and dense scenarios. Furthermore, by changing the initial and target positions, different path planning scenarios can be tested, as shown in Fig.4d.
@@ IV-B2 Results in dynamic environment
Real-time simulation was also performed on dynamic environments involving the continuous movement of one ground obstacle and one obstacle that was deflected by both the drone and mobile robot. Fig. 5 shows the tests that were performed in two main cases with sparse (Fig. 5a) and dense (Fig. 5b) environments. The change of path can be observed in the region where the movement of the obstacle takes place. The same behavior is observed in a cluttered environment, shown in Fig.5c and Fig.5d. The results revealed the adaptable behavior of a heterogeneous swarm under multiple environmental conditions.
Figure 5: Simulation examples of dynamic obstacle avoidance in sparse (a, b) and dense (c, d) environments. Dashed purple lines are the obstacle trajectories. Black dots are the obstacles for both robots, and green dots are the ground obstacles deflected only by the mobile robot. Gray and light green-colored circles show the safe deflection regions for both agents.
@@ IV-B5 Time efficiency in task completion for multiple agents
To check the robustness and system adaptability under varying conditions, we conducted experiments in two different environments, shown in Fig. 4c, and Fig. 4d, with varying starting and goal positions, both in a static and dynamic environment. Table II shows how the trajectory length and mission time change under varying environmental conditions.
@@ IV-B6 Success Rate
$\text{Success Rate}=\left(\frac{\text{Successful Experiments}}{\text{Total Experiments}}\right)\times 100,$ (4)
@@ IV-C Comparison with Conflict-Based Search (CBS) Algorithm
a 0.73 0.25 0.75 0.38
b 0.72 0.25 0.75 0.37
With similar environmental configurations, the CBS trajectory in Case 1, where there are fewer obstacles, shows that agents tend to move around obstacles, resulting in a longer path. In contrast, APF agents navigate within obstacles, leading to more efficient paths. However, in Case 2, where the environment is more densely packed with obstacles, CBS agents ignore the inner regions and move around obstacles to complete the trajectory. On the other hand, APF agents move between obstacles, encountering multiple deflections, which increase their path length. Despite this, APF excels in minimum obstacle distance; agents maintain a very close distance to obstacles while still avoiding collisions. In comparison, CBS agents maintain a minimum distance of 72 cm, reflecting a more cautious but less flexible path.
@@ IV-D Conclusion and Future Work
The research presents a heterogeneous swarm system comprising a drone and a mobile robot and designed to collaborate in dynamic environments for logistics and warehouse management tasks. This research works on a leader-follower approach where a drone, being a leader, uses an APF path planner to navigate to goal positions while avoiding obstacles. Moreover, the mobile robot follows the drone using impedance links that connect the drone with the mobile robot. Moreover, the mobile robot develops additional impedance links with the ground obstacles to ensure smooth navigation toward the goal.
By combining the two agents, the system proves to have the complementary strengths of both a drone and a mobile robot. Multiple experiments were performed on this system, achieving a success rate of 90%. Additionally, a deviation of 45 cm was observed in the mobile robot’s path near the ground obstacles region, indicating that the robot is effectively avoiding them. The current approach was later compared with the Conflict-Based Search (CBS) algorithm, which reveals that our approach enables agents to safely yet optimally navigate within 25 cm of obstacles, whereas CBS maintains a minimum distance of 73 cm.
@@
Acknowledgements
Research reported in this publication was financially supported by the RSF-DST grant No. 24-41-02039.